# DAWN: Designing Distributed Agents in a Worldwide Network

### v2

@@ IV Application of DAWN
Every step of the hiring process allocated to the agentic node involves the Principal Agent and Gateway Agent working together, showcasing the autonomous capabilities of DAWN’s framework. On the other hand, the No-LLM node demonstrates a more deterministic approach where a human takes over, primarily conducted through deterministic steps. This step includes offer negotiation and background checking, which involve candidate interaction and highly confidential and sensitive data that must be kept secure. By using the No-LLM mode, the DAWN framework allows for the integration of full determinism in the workflow when necessary. This HR use case was implemented with the listed resources to demonstrate the functionality of our framework, and leverages both expert knowledge and the autonomy of using LLM agents in applications requiring sophisticated task execution.
V Conclusion and Future Directions
@@
The DAWN platform revolutionizes agentic applications by establishing a robust foundation for agent collaboration and communication. This framework offers a comprehensive strategy for integrating LLM-based agents into diverse applications. Its modular design, featuring key components like the Principal Agent and Gateway Agent, ensures that resources are optimally utilized, enabling seamless workflow management and global agent cooperation.
The DAWN platform’s contributions—hierarchical architecture engendering flexibility, scalability, interoperability, and stringent safety and compliance measures—position it as a transformative solution for small and large applications. By addressing the challenges of robustness and reliability in agent-based systems, DAWN paves the way for advanced utilization across various industries, from customer support to cybersecurity. The continued development and refinement of the DAWN framework is expected to push the boundaries of what is possible with LLM-based agents, fostering an interconnected ecosystem where collaboration and innovation thrive.
@@
The evaluation of complex agentic workflows remains a dynamic field of research, with considerable focus on developing benchmarks and metrics for multi-agent applications. The lack of standardized benchmarks poses challenges, as does the assessment of individual network components in isolation—such as the Principal agent—due to the inherent complexity and non-determinism in multi-agent networks. Establishing clear metrics and success measures for multi-agent systems is essential, and DAWN’s architecture encourages further research and refinement in this field, potentially setting standards for agentic networks of the future.
VI Acknowledgment
This research is supported by Outshift at Cisco Systems. We extend our profound thanks to Reinaldo for his diligent review of our paper and for providing expertise that greatly contributed to enhancing the second version of the DAWN paper. We also express our sincere appreciation to Nihar Dandekar for sharing insights on potential solutions for agentic frameworks

### v3

@@ IV Application of DAWN
Each of these agentic steps executed in seconds with most of the elapsed time spent in LLM interactions. When run end-to-end, the runs varied from eight to 20 seconds using GPT-4o with all of the variance in elapsed time attributable to LLM inference times. Every step of the hiring process allocated to an agentic node involved the Principal Agent and Gateway Agent working together, showcasing the copilot and autonomous capabilities of this DAWN implementation. In this HR application, the task graph (agentic workflow) awaits human interaction after each agentic step - a requirement of the designer. Once completed the workflow progresses to the next step automatically. The No-LLM workflow step demonstrates the inclusion of a traditional non-AI business process where the hiring manager takes over. This step includes salary estimation, negotiations, and candidate background checks which involve interactions between multiple parties and highly confidential and sensitive data that must be kept secure. Once complete, the workflow resumes where it left off, but requires specific prompting. (It also handles concurrency when multiple such requests are in flight, and prompting for resumption requires additional data to pick up the pending actions.) The No-LLM mode illustrates how the DAWN architecture allows for the integration of human-in-the-loop and determinism within an application’s agentic workflow when necessary. This HR use case was implemented with the listed resources to demonstrate the functionality of this framework, and leverages both expert knowledge and the autonomy of using LLM agents in applications requiring sophisticated task execution.
For each step of the workflow, the Principal Agent communicated back and forth with the connected Gateway Agents for further action, ensuring the plan is regularly evaluated and the user interface (and user) is updated with operational status. This requires that state is continuously exchanged between the Principal and the Gateway Agents for error management, telemetry data collection, and status updates. This continuous communication is essential to support iterative and traceable problem-solving, and ensure transparency and accountability throughout the entire workflow execution process, thus making complex task management more effective. Yet another implementation using the DAWN architecture is described in this Cisco Tech Blog [69]. This implementation also deploys one Principal Agent and two Gateway Agents but this time using three different agentic frameworks, LangGraph, Swarm, and AutoGen (or AG2), for agent creation, deployment, and execution. Execution times tracked what was measured in the HR use case. Together these two implementations demonstrate the flexibility and viability of the DAWN architecture.
V Conclusion and Future Direction
@@
The DAWN platform revolutionizes agentic applications by establishing a robust foundation for agent collaboration and communication. This framework offers a comprehensive strategy for integrating LLM-based agents into diverse applications, both new and existing. Its modular design, featuring key services like the Principal Agent and Gateway Agents, ensures that distributed resources are optimally utilized, enabling seamless workflow management and global agent cooperation across different organizations. The architecture also enables small and large participants to collaborate across the lifecyle of agentic systems.
The DAWN platform’s contributions—hierarchical architecture engendering flexibility, scalability, interoperability, safety, and compliance—position it as a transformative solution for applications of varying complexity. It shows how agents may be registered, discovered, connected and threaded into workflows to meet the needs of agentic software development. DAWN also provides control points to address the safety, security, privacy, and compliance requirements of agentic software development within both the Principal Agent and Gateway Agents. The Principal Agent is wholly owned and operated by the organization that develops it and therefore its make up falls entirely within its purview. Gateway Agents, on the other hand, may be owned operated by subsidiaries, partners, and even competitors. The DAWN architecture therefore provides for maximum flexibility while also allowing for requisite compliance considerations.
@@
The evaluation of complex agentic workflows remains a dynamic field of research, with considerable focus on developing benchmarks and metrics for multi-agent applications. The lack of standardized benchmarks poses challenges, as does the assessment of individual network components in isolation – such as the Principal Agent – due to the inherent complexity and non-determinism in multi-agent networks. Establishing clear metrics and success measures for multi-agent systems is essential, and DAWN’s architecture encourages further research and refinement in this field, potentially setting standards for agentic networks of the future.
VI Acknowledgments
This research is supported by Outshift AI, the innovation engine of Cisco Systems, Inc. We extend our profound thanks to Reinaldo Penno, Distinguished Engineer, Outshift, for his diligent review of our paper. We also express our sincere appreciation to Nihar Dandekar, Director, Outshift, for sharing insights on potential solutions for agentic frameworks.